# The metric theory of small gaps for a sequence of real numbers

### v1

@@ 5. Proof of Theorem 13
See [kuipers1974uniform, Theorem 4.1] for a proof. ∎
@@
For each positive integer $N$, let $T(N)\coloneqq\#\{d\mid N:d\geq\log N\}$, and we enumerate the divisors as
$\{d\mid N:d\geq\log N\}=\{d_{1}^{N},d_{2}^{N},\dots,d_{T(N)}^{N}\},$
@@
Now, for each $d\mid k_{2}$ satisfying $d\geq\log k_{2}$, we have $d\geq k_{1}$, that is, $d>d_{i}^{k_{1}}$ for each $i=1,\ldots,T(k_{1})$. Thus $d_{j}^{k_{2}}>d_{i}^{k_{1}}$ for all $i\in\{1,\ldots,T(k_{1})\}$ and $j\in\{1,\ldots,T(k_{2})\}$.
@@
By construction, we have $d_{i}^{k_{n-1}}<d_{j}^{k_{n}}$ for all $i\in\{1,\ldots,T(k_{n-1})\}$ and $j\in\{1,\ldots,T(k_{n})\}$. Also, for fixed $N$, $d_{r}^{N}<d_{s}^{N}$ for $1\leq r<s\leq T(N)$. Therefore,
@@
  1. (1)
    This is a standard number theory result.
  2. (2)
    Note that $\varphi(d)\leq d$ for all $d\in\mathbb{N}$. Therefore,
@@
    The last inequality holds for all sufficiently large $k$, which concludes the proof.
@@
for each $N\in\mathbb{N}$. This shows that $A$ is an infinite set of positive integers, which is also countable.
@@
Since $(Z_{n})_{n\geq 1}$ is a sequence of distinct integers, Weyl’s equidistribution theorem (see Lemma 59) implies that for almost every $\alpha$, the sequence $(\alpha Z_{n})_{n\geq 1}$ is uniformly distributed modulo 1. Hence, for any $\delta>0$,
@@
Since a sequence of distinct integers has an increasing subsequence, by passing to one such subsequence $(Z_{l(n)})_{n\geq 1}$ of $(Z_{n})_{n\geq 1}$ and applying Lemma 58, we find that $(\alpha Z_{l(n)})_{n\geq 1}$ is uniformly distributed for almost all $\alpha\in[0,1)$. Hence, Lemma 58 can indeed be used in place of Lemma 59 in the proof of Lemma 63, as noted in [ABM].
@@
Case 1: Let $\limsup_{k\to\infty}\eta(k)=u^{*}>0$. We will show that the series (8) diverges and that $\lambda(\mathcal{A})=1$. By Lemma 63, the inequality $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$ has infinitely many solutions for almost all $\alpha\in[0,1]$. Consequently, $\lambda(\mathcal{A})=1$.
@@
for sufficiently large $k\in A$, where $\mathcal{N}(k)<\infty$ ensures that the supremum is finite (The condition $\mathcal{N}(k)<\infty$ ensures that we are not taking a supremum over an empty set, in which case it would be 0, by convention). Since $A$ is an infinite set, we use (183) and Corollary 61 to deduce that (8) is indeed divergent in this case.
@@
and we also set $S_{k}=\emptyset$ for all $k\notin A$. Let
$\mathcal{A}=\{\alpha\in[0,1]:\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)\text{ for infinitely many }N\},$
defined as in the statement of the theorem, and
@@
converges or diverges, respectively. Note that $\mathcal{N}(k)=\infty$ for $k\notin A$, and the supremum over the empty set is zero by convention. Since $\eta(\ell)\geq 0$, both
(188) $\displaystyle\sup_{b\geq 1}\left\{\frac{\psi(bk)}{bk}\right\}\text{ and }\displaystyle\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right\}$
coincide with $\sup_{\begin{subarray}{c}b\geq 1\\ bk\in A\end{subarray}}\left(\frac{\sup_{l\geq\mathcal{N}(bk)}\eta(l)}{bk}\right)$ for all $k\in\mathbb{N}$. Therefore, the two series in (187) are equal.
To show $\mathcal{A}=\mathcal{B}$, assume first that $\alpha\in\mathcal{A}$. Then, there are infinitely many values of $N,m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. Set $k=\lfloor|a_{m}-a_{n}|\rfloor$. Since $m,n\leq N$, we have $\mathcal{N}(k)\leq N$. Thus, we have $\psi(k)=\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)\geq\eta(N)$. Consequently, there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, that is, $\|k\alpha\|\leq\psi(k)$, and thus $\alpha\in S_{k}$. Since $\eta(N)\rightarrow 0$, a particular difference $k=\lfloor|a_{m}-a_{n}|\rfloor$ can only generate finitely many values $N$ (together with $m,n\leq N)$ such that $\|\lfloor|a_{m}-a_{n}|\rfloor\alpha\|\leq\eta(N)$. Since each $k$ corresponds to finitely many $N$, and there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, the choices for $k$ satisfying $\|k\alpha\|\leq\eta(N)$ must be infinite. In other words, there are infinitely many distinct values of $k$ such that $\alpha\in S_{k}$, and hence $\mathcal{A}\subseteq\mathcal{B}$.
We now assume that $\alpha\notin\mathcal{A}$. Then, there are only finitely many $N$ such that $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$, that is, there are only finitely many $N$ and $m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. This implies that there are only finitely many $k,N\in\mathbb{N}$ such that $k\in\lfloor(A_{N}-A_{N})^{+}\rfloor$ and $\|k\alpha\|\leq\eta(N)$. Since $\eta(N)\rightarrow 0$, there are only finitely many $k$, and for each such $k$, there are only finitely many $\ell\geq\mathcal{N}(k)$ satisfying $\|k\alpha\|\leq\eta(l)$. Therefore, there are only finitely many $k$ for which $\|k\alpha\|\leq\sup_{l\geq\mathcal{N}(k)}\eta(l)$, that is, $\|k\alpha\|\leq\psi(k)$. Thus, there are only finitely many $k$ such that $\alpha\in S_{k}$.
@@ Acknowledgements
The author is grateful to Christoph Aistleitner and Sudhir Pujahari for thoughtful inputs and correspondence. The author is supported by a doctoral fellowship from the National Board for Higher Mathematics and the second named author was supported by the Department of Atomic Energy (DAE project number RIN-4001).

### v2

@@ 5. Proof of Theorem 13
See [kuipers1974uniform, Theorem 4.1]. ∎
@@
For each positive integer $N$, let $T(N)\coloneqq\#\{d:d\mid N,\,d\geq\log N\}$, and we enumerate the divisors as
$\{d:d\mid N,\,d\geq\log N\}=\{d_{1}^{N},d_{2}^{N},\dots,d_{T(N)}^{N}\},$
@@
Now, for each $d\mid k_{2}$ satisfying $d\geq\log k_{2}$, we have $d\geq k_{1}$, that is, $d>d_{i}^{k_{1}}$ for each $i=1,\ldots,T(k_{1})$. Thus, $d_{j}^{k_{2}}>d_{i}^{k_{1}}$ for all $i\in\{1,\ldots,T(k_{1})\}$ and $j\in\{1,\ldots,T(k_{2})\}$.
@@
By construction, we have $d_{i}^{k_{m}}<d_{j}^{k_{n}}$ for all $m<n$, $i\in\{1,\ldots,T(k_{m})\}$, and $j\in\{1,\ldots,T(k_{n})\}$. Also, for fixed $N$, $d_{r}^{N}<d_{s}^{N}$ for $1\leq r<s\leq T(N)$. Therefore,
@@
We only prove $(2)$, as $(1)$ is a standard fact in number theory. Since $\varphi(d)\leq d$ for all $d\in\mathbb{N}$, we have
@@
The last inequality holds for all sufficiently large $k$, which concludes the proof. ∎
@@
for each $N\in\mathbb{N}$. This shows that $A$ is an infinite set of positive integers.
@@
Since $(Z_{n})_{n\geq 1}$ is a sequence of distinct integers, Weyl’s equidistribution theorem (see Lemma 58) implies that for almost every $\alpha$, the sequence $(\alpha Z_{n})_{n\geq 1}$ is uniformly distributed mod $1$. Hence, for any $\delta>0$,
@@
Since a sequence of distinct integers has an increasing subsequence, by passing to one such subsequence $(Z_{l(n)})_{n\geq 1}$ of $(Z_{n})_{n\geq 1}$ and applying Lemma 57, we find that $(\alpha Z_{l(n)})_{n\geq 1}$ is uniformly distributed for almost all $\alpha\in[0,1)$. Hence, Lemma 57 can be used in place of Lemma 58 in the proof of Lemma 62, as noted in [ABM].
@@
Case 1: Let $\limsup_{k\to\infty}\eta(k)=u^{*}>0$. We will show that the series (8) diverges and that $\lambda(\mathcal{A})=1$. By Lemma 62, inequality $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$ has infinitely many solutions for almost all $\alpha\in[0,1]$. Consequently, $\lambda(\mathcal{A})=1$.
@@
for sufficiently large $k\in A$, where $\mathcal{N}(k)<\infty$ ensures that the supremum is finite (condition $\mathcal{N}(k)<\infty$ ensures that we are not taking a supremum over an empty set, in which case it would be 0, by convention). Since $A$ is an infinite set, we use (180) and Corollary 60 to deduce that (8) is indeed divergent in this case.
@@
and set $S_{k}=\emptyset$ for $k\notin A$. Let
$\mathcal{A}=\{\alpha\in[0,1]:\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)\text{ for infinitely many }N\}$
be defined as in the statement of the theorem, and
@@
converges or diverges, respectively. Note that $\mathcal{N}(k)=\infty$ for $k\notin A$, and the supremum over the empty set is zero by convention. Since $\eta(\ell)\geq 0$, both expressions
(185) $\displaystyle\sup_{b\geq 1}\left\{\frac{\psi(bk)}{bk}\right\}\quad\text{and}\quad\displaystyle\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right\}$
coincide with $\sup_{\begin{subarray}{c}b\geq 1\\ bk\in A\end{subarray}}\left(\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right)$ for all $k\in\mathbb{N}$. Therefore, the two series in (184) are equal.
To show $\mathcal{A}=\mathcal{B}$, first assume that $\alpha\in\mathcal{A}$. Then, there are infinitely many values of $N,m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. Set $k=\lfloor|a_{m}-a_{n}|\rfloor$. Since $m,n\leq N$, we have $\mathcal{N}(k)\leq N$. Thus, we have $\psi(k)=\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)\geq\eta(N)$. Consequently, there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, that is, $\|k\alpha\|\leq\psi(k)$, and thus $\alpha\in S_{k}$. Since $\eta(N)\rightarrow 0$, a particular difference $k=\lfloor|a_{m}-a_{n}|\rfloor$ can generate only finitely many values $N$ (together with $m,n\leq N)$ such that $\|\lfloor|a_{m}-a_{n}|\rfloor\alpha\|\leq\eta(N)$. Since each $k$ corresponds to finitely many $N$, and there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, the choices for $k$ satisfying $\|k\alpha\|\leq\eta(N)$ must be infinite. In other words, there are infinitely many distinct values of $k$ such that $\alpha\in S_{k}$, and therefore $\mathcal{A}\subseteq\mathcal{B}$.
We now assume that $\alpha\notin\mathcal{A}$. Then, there are only finitely many $N$ such that $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$, that is, there are only finitely many $N$ and $m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. This implies that there are only finitely many $k,N\in\mathbb{N}$ such that $k\in\lfloor(A_{N}-A_{N})^{+}\rfloor$ and $\|k\alpha\|\leq\eta(N)$. Since $\eta(N)\rightarrow 0$, there are only finitely many $k$, and for each such $k$, there are only finitely many $\ell\geq\mathcal{N}(k)$ satisfying $\|k\alpha\|\leq\eta(l)$. Therefore, there are only finitely many $k$ for which $\|k\alpha\|\leq\sup_{l\geq\mathcal{N}(k)}\eta(l)$, that is, $\|k\alpha\|\leq\psi(k)$. Thus, there are only finitely many $k$ such that $\alpha\in S_{k}$, which shows that $\alpha\notin\mathcal{B}$.
@@ Acknowledgements
The author is grateful to Christoph Aistleitner and Sudhir Pujahari for thoughtful inputs and correspondence. The author is supported by the Department of Atomic Energy (DAE project number RIN-4001).